IMPORTANT STEP: For each experiment, Fixed Phen and Fixed Fe, write the Beers Law equation for each sloped series trend line. Do not use the y-intercept term, and do not use y or x as the variables. Use the chemical variables A for absorbance and the reactant concentration symbolism, []. For example, if the trendline equation for the varying mmoles of iron (II) sulfate is y-1070x, then the equation you would use is: A-1070 [Fe] A 1062 el If the trendline equation for the varying mmoles of phenanthroline is y-341.03x, then the equation you would use is: A-341 [Ph] A-294.8 Using YOUR equations, solve the two equations simultaneously. One way to do this is to SUBSTITUTE 341 [Ph] into the other equation and then solve for the reaction mole ratio of the two species. This ratio should be close to a whole number. If not, review how you split your data series to try and improve your R value. This graphical and mathematical process is the answer to Question #3 below, and the number will be the value of x in the following chemical equation:
Answer #1

You have 2 equations for absorbance

A = 1067.6 [Fe] and A = 294.8 [Ph], from the statement you want to relate this 2 equations like this

1067.6 [Fe] = 294.8 [Ph], you want to get a mole ratio or these 2 species so if we rearrange the numbers:

1067.6 [Fe] / 294.8 = [Ph]

1067.6 / 294.8 = [Ph] / [Fe]

3.6214 = [PH] / [Fe]

if you want it inverse

[Fe] / [PH] = 1 / 3.6214 = 0.276

*Please rate the answer if you like it =)